1-8. How much could a company afford to spend on an HVAC system that would bring monthly savings of S1000 over the entire 12-year life of the equipment? The company uses an annual interest rate of 12 percent in making investment projections.
Answer #1

Monthly savings = 1000
Rate per month = 12%/12 = 1%
Number of Periods = 12*12 = 144
PV = 1000*(1-(1+1%)-144)/1% = 76,137.16

The company can afford to spend 76137.16 on HVAC

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well